OpenStudy (anonymous):

Factor 4a2 - 25. (2a + 5)(2a + 5) (2a - 5)(2a - 5) (2a - 5)(2a + 5)

OpenStudy (anonymous):

a?

OpenStudy (solomonzelman):

\[4a^2-25\]

OpenStudy (kc_kennylau):

Use the identity of the difference of squares: \((a+b)(a-b)=a^2-b^2\)

OpenStudy (anonymous):

and yes @SolomonZelman

OpenStudy (kc_kennylau):

and no @dirtydan667 not a

OpenStudy (solomonzelman):

Opps, sorry

OpenStudy (anonymous):

c?

OpenStudy (solomonzelman):

\[(~~(2a)^2-5^2~)\]

OpenStudy (anonymous):

meant b

OpenStudy (solomonzelman):

use @kc_kennylau 's identity.

OpenStudy (kc_kennylau):

Since it's the second time u ask this kind of questions, I'll assume that you do not know how to do other questions of this type. There are five identities used for this kind of question. 1. \(\large(a+b)^2=a^2+2ab+b^2\) 2. \(\large(a-b)^2=a^2-2ab+b^2\) 3. \(\large a^2-b^2=(a+b)(a-b)\) 4. \(\large a^3+b^3=(a+b)(a^2-ab+b^2)\) 5. \(\large a^3-b^3=(a-b)(a^2+ab+b^2)\) You just need to know this five identities by heart and you shall know how to do quesitons of this type.

OpenStudy (kc_kennylau):

I suggest you to consult your teacher for further inquiries.